OpenStudy (anonymous):

he traffic cone is 19 inches tall & has a radius of 5 inches. Find the surface area.

OpenStudy (anonymous):

surface area=pi*r*l+pi*r*r l=square root of h*h+r*r putting values if get l=square root of 386 l=19.647 putting value for surface area= SA = pi*5*5+pi*5*19.647 SA=387.15